# Linear Contextual Bandits with Hybrid Payoff: Revisited

### v1

@@ Abstract
We study the Linear Contextual Bandit (LinearCB) problem in the hybrid reward setting. In this setting, every arm’s reward model contains arm specific parameters in addition to parameters shared across the reward models of all the arms. We can easily reduce this setting to two closely related settings; (a) Shared - no arm specific parameters, and (b) Disjoint - only arm specific parameters, enabling the application of two popular state of the art algorithms - LinUCB and DisLinUCB (proposed as Algorithm $1$ in (Li et al., 2010)). When the arm features are stochastic and satisfy a popular diversity condition, we provide new regret analyses for both LinUCB and DisLinUCB that significantly improves upon the known regret guarantees of these algorithms. Our novel analysis critically exploits the structure of the hybrid rewards and diversity of the arm features. Along with proving these new guarantees, we introduce a new algorithm HyLinUCB that crucially modifies LinUCB (using a new exploration coefficient) to account for sparsity in the hybrid setting. Under the same diversity assumptions, we prove that at the end of $T$ rounds, HyLinUCB also incurs only $\tilde{O}(\sqrt{T})$ regret. We perform extensive experiments on synthetic and real-world datasets demonstrating strong empirical performance of HyLinUCB. When the number of arm specific parameters is much larger than the number of shared parameters, we observe that DisLinUCB incurs the lowest regret. In this case, regret of HyLinUCB is the second best and it is extremely competitive to DisLinUCB. In all other situations, including our real-world dataset, HyLinUCB has significantly lower regret than LinUCB, DisLinUCB and other state of the art baselines we considered. We also empirically observe that the regret of HyLinUCB grows much slower with the number of arms $K$, compared to baselines, making it suitable even for very large action spaces.
@@ 1.3 Related Work
As highlighted in Section 1.2, the LinearCB problem in the hybrid setting can be reduced to the shared setting, enabling the application of all LinearCB algorithms designed for the shared setting. Over the last couple of decades, there has been a substantial progress on LinearCB for the shared setting. Many state of the art algorithms follow the optimism in the face of uncertainty approach (Lai and Robbins, 1985), and provide regret analysis with near optimal guarantees (Auer, 2003; Dani et al., 2008; Rusmevichientong and Tsitsiklis, 2008; Chu et al., 2011; Abbasi-Yadkori et al., 2011). For general stochastic arm features, the best known regret upper bound (independent of number of arms $K$) provided in these works is $\tilde{O}(d\sqrt{T})$ and the best known lower bound (Chu et al., 2011) is $\tilde{\Omega}(\sqrt{dT})$, where $d$ is the dimensionality of the arm features and $T$ is the number of rounds. The LinUCB algorithm (Li et al., 2010) attains this upper bound ($\tilde{O}(d\sqrt{T})$) Lattimore and Szepesvári (2020) and is also a popular choice for applications such as news recommendation (Li et al., 2010) due to it’s strong empirical performance. Moreover, it was also adapted to the hybrid setting in Li et al. (2010). As a result we choose it as one of our main candidate algorithms to study in this paper. The popular SupLinUCB (Chu et al., 2011) algorithm improves the dependence on $d$ and guarantees $\tilde{O}(\sqrt{dT}\log^{3/2}{K})$ regret where $K$ is the number of actions. A recent variant of the SupLinUCB algorithm (Li et al., 2019) improves this further to $\tilde{O}(\sqrt{dT}\log{K})$. While the dependence on $d$ improves, it is traded off with a logarithmic dependence on $K$. It has also been noted (Kim et al., 2023) that despite strong regret guarantees, SupLinUCB and its variants (Li et al., 2017; 2019) explore excessively and perform worse in practice due to computational inefficiency. Recently, Kim et al. (2023) developed a new algorithm HyRan which does not depend on $K$ in the leading terms and attains a regret guarantee of $\tilde{O}(\sqrt{dT})$ (improving significantly over SupLinUCB both theoretically and empirically). While they prove this under an additional assumption on stochasticity of the arm features, they also provide a matching lower bound showing tightness of their guarantee. Their experiments validate their strong upper bound by demonstrating superior performance compared to many state of the art LinearCB baselines for a large variety of problem instances. Due to their strong guarantee and empirical performance, we compare our algorithms with HyRan in Section 4. Another family of algorithms (based on randomized exploration) referred to as Thompson Sampling (Thompson, 1933) based methods are an active area of research. Linear contextual version of this method was recently developed (Agrawal and Goyal, 2013; Abeille and Lazaric, 2017) and best known regret guarantees are either $\tilde{O}(d\sqrt{T\log{K}})$ or $\tilde{O}(d^{3/2}\sqrt{T})$. Both guarantees are worse than that of LinUCB. Even though Thompson Sampling does well practically, the HyRan algorithm from Kim et al. (2023) was shown to be superior to them for all problem instances considered in Kim et al. (2023). Since we already compare with HyRan in our experiments in Section 4, we do not compare to Thompson Sampling based methods.
@@ 3.1 LinUCB and DisLinUCB
At the end of $T$ rounds, the regret of LinUCB (Algorithm 1 with $\lambda=1$, $\gamma=S\sqrt{K}+\sqrt{2(d_{1}+d_{2}K)\log(T/\delta)}$) under Assumptions 1 and 2 is upper bounded by $C\sqrt{(d_{1}+d_{2}K)T\log(T/\delta)}$ with probability at least $1-4\delta$. Here, $C>0$ is a universal constant and $T$ is assumed to be $\tilde{\Omega}(K^{4})$.
@@
5:     $i_{t}=\mathop{\mathrm{argmax}}_{i\in[K]}\langle{\overline{x}_{i}},{\phi_{i}}\rangle+\gamma\left\lVert\overline{x}_{i}\right\rVert_{\mathbf{V}_{i}}$
@@
At the end of $T$ rounds, the regret of DisLinUCB (Algorithm 2 with $\lambda=1$, $\gamma=2\sqrt{S}+\sqrt{2(d_{1}+d_{2})\log(KT/\delta)}$) under Assumptions 1 and 2 is upper bounded by $C\sqrt{(d_{1}+d_{2})T\log(KT/\delta)}$ with probability at least $1-4\delta$. Here, $C>0$ is a universal constant and $T$ is assumed to be $\tilde{\Omega}(1)$.
@@ A.1 Proof of Theorem 3
At the end of $T$ rounds, the regret of LinUCB (Algorithm 1 with $\lambda=1$, $\gamma=2S\sqrt{K}+\sqrt{2(d_{1}+d_{2}K)\log(T/\delta)}$) under Assumptions 1, 2 is upper bounded by $C\sqrt{(d_{1}+d_{2}K)T\log(T/\delta)}$ with probability at least $1-4\delta$. Here, $C>0$ is a universal constant and $T$ is assumed to be $\tilde{\Omega}(K^{4})$.
@@ Appendix C Regret Analysis of DisLinUCB
At the end of $T$ rounds, the regret of DisLinUCB (Algorithm 2 with $\lambda=1$, $\gamma=2\sqrt{S}+\sqrt{2(d_{1}+d_{2})\log(KT/\delta)}$) under Assumptions 1, 2 is upper bounded by $C\sqrt{(d_{1}+d_{2})T\log(KT/\delta)}$ with probability at least $1-4\delta$. Here, $C>0$ is a universal constant and $T$ is assumed to be $\tilde{\Omega}(1)$.

### v2

@@ Abstract
We study the Linear Contextual Bandit (LinearCB) problem in the hybrid reward setting. In this setting, every arm’s reward model contains arm specific parameters in addition to parameters shared across the reward models of all the arms. We can easily reduce this setting to two closely related settings; (a) Shared - no arm specific parameters, and (b) Disjoint - only arm specific parameters, enabling the application of two popular state of the art algorithms - LinUCB and DisLinUCB (proposed as Algorithm $1$ in Li et al. 2010). When the arm features are stochastic and satisfy a popular diversity condition, we provide new regret analyses for both LinUCB and DisLinUCB that significantly improves upon the known regret guarantees of these algorithms. Our novel analysis critically exploits the structure of the hybrid rewards and diversity of the arm features. Along with proving these new guarantees, we introduce a new algorithm HyLinUCB that crucially modifies LinUCB (using a new exploration coefficient) to account for sparsity in the hybrid setting. Under the same diversity assumptions, we prove that at the end of $T$ rounds, HyLinUCB also incurs only $\tilde{O}(\sqrt{T})$ regret. We perform extensive experiments on synthetic and real-world datasets demonstrating strong empirical performance of HyLinUCB. When the number of arm specific parameters is much larger than the number of shared parameters, we observe that DisLinUCB incurs the lowest regret. In this case, regret of HyLinUCB is the second best and it is extremely competitive to DisLinUCB. In all other situations, including our real-world dataset, HyLinUCB has significantly lower regret than LinUCB, DisLinUCB and other state of the art baselines we considered. We also empirically observe that the regret of HyLinUCB grows much slower with the number of arms $K$, compared to baselines, making it suitable even for very large action spaces.
@@ 1.3 Related Work
As highlighted in Section 1.2, the LinearCB problem in the hybrid setting can be reduced to the shared setting, enabling the application of all LinearCB algorithms designed for the shared setting. Over the last couple of decades, there has been a substantial progress on LinearCB for the shared setting. Many state of the art algorithms follow the optimism in the face of uncertainty approach (Lai and Robbins, 1985), and provide regret analysis with near optimal guarantees (Auer, 2003; Dani et al., 2008; Rusmevichientong and Tsitsiklis, 2008; Chu et al., 2011; Abbasi-Yadkori et al., 2011). For general stochastic arm features, the best known regret upper bound (independent of number of arms $K$) provided in these works is $\tilde{O}(d\sqrt{T})$ and the best known lower bound (Chu et al., 2011) is $\tilde{\Omega}(d\sqrt{T})$, where $d$ is the dimensionality of the arm features and $T$ is the number of rounds. The LinUCB algorithm (Li et al., 2010) attains this upper bound ($\tilde{O}(d\sqrt{T})$) Lattimore and Szepesvári (2020) and is also a popular choice for applications such as news recommendation (Li et al., 2010) due to it’s strong empirical performance. Moreover, it was also adapted to the hybrid setting in Li et al. (2010). As a result we choose it as one of our main candidate algorithms to study in this paper. The popular SupLinUCB (Chu et al., 2011) algorithm improves the dependence on $d$ and guarantees $\tilde{O}(\sqrt{dT}\log^{3/2}{K})$ regret where $K$ is the number of actions. A recent variant of the SupLinUCB algorithm (Li et al., 2019) improves this further to $\tilde{O}(\sqrt{dT}\log{K})$. While the dependence on $d$ improves, it is traded off with a logarithmic dependence on $K$. It has also been noted (Kim et al., 2023) that despite strong regret guarantees, SupLinUCB and its variants (Li et al., 2017; 2019) explore excessively and perform worse in practice due to computational inefficiency. Recently, Kim et al. (2023) developed a new algorithm HyRan which does not depend on $K$ in the leading terms and attains a regret guarantee of $\tilde{O}(\sqrt{dT})$ (improving significantly over SupLinUCB both theoretically and empirically). While they prove this under an additional assumption on stochasticity of the arm features, they also provide a matching lower bound showing tightness of their guarantee. Their experiments validate their strong upper bound by demonstrating superior performance compared to many state of the art LinearCB baselines for a large variety of problem instances. Due to their strong guarantee and empirical performance, we compare our algorithms with HyRan in Section 4. Another family of algorithms (based on randomized exploration) referred to as Thompson Sampling (Thompson, 1933) based methods are an active area of research. Linear contextual version of this method was recently developed (Agrawal and Goyal, 2013; Abeille and Lazaric, 2017) and best known regret guarantees are either $\tilde{O}(d\sqrt{T\log{K}})$ or $\tilde{O}(d^{3/2}\sqrt{T})$. Both guarantees are worse than that of LinUCB. Even though Thompson Sampling does well practically, the HyRan algorithm from Kim et al. (2023) was shown to be superior to them for all problem instances considered in Kim et al. (2023). Since we already compare with HyRan in our experiments in Section 4, we do not compare to Thompson Sampling based methods.
@@ 3.1 LinUCB and DisLinUCB
At the end of $T$ rounds, the regret of LinUCB (Algorithm 1 with $\lambda=1$, $\gamma=S\sqrt{K}+\sqrt{2(d_{1}+d_{2}K)\log(T/\delta)}$) under Assumptions 1 and 2 is upper bounded by $C\sqrt{(d_{1}+d_{2}K)KT\log(T/\delta)}$ with probability at least $1-4\delta$. Here, $C>0$ is a universal constant and $T$ is assumed to be $\tilde{\Omega}(K^{4})$.
@@
5:     $i_{t}=\mathop{\mathrm{argmax}}_{i\in[K]}\langle{\overline{x}_{i}},{\phi_{i}}\rangle+\gamma\left\lVert\overline{x}_{i}\right\rVert_{\mathbf{V}_{i}^{-1}}$
@@
At the end of $T$ rounds, the regret of DisLinUCB (Algorithm 2 with $\lambda=1$, $\gamma=2\sqrt{S}+\sqrt{2(d_{1}+d_{2})\log(KT/\delta)}$) under Assumptions 1 and 2 is upper bounded by $C\sqrt{(d_{1}+d_{2})KT\log(KT/\delta)}$ with probability at least $1-4\delta$. Here, $C>0$ is a universal constant and $T$ is assumed to be $\tilde{\Omega}(1)$.
@@ A.1 Proof of Theorem 3
At the end of $T$ rounds, the regret of LinUCB (Algorithm 1 with $\lambda=1$, $\gamma=2S\sqrt{K}+\sqrt{2(d_{1}+d_{2}K)\log(T/\delta)}$) under Assumptions 1, 2 is upper bounded by $C\sqrt{(d_{1}+d_{2}K)KT\log(T/\delta)}$ with probability at least $1-4\delta$. Here, $C>0$ is a universal constant and $T$ is assumed to be $\tilde{\Omega}(K^{4})$.
@@ Appendix C Regret Analysis of DisLinUCB
At the end of $T$ rounds, the regret of DisLinUCB (Algorithm 2 with $\lambda=1$, $\gamma=2\sqrt{S}+\sqrt{2(d_{1}+d_{2})\log(KT/\delta)}$) under Assumptions 1, 2 is upper bounded by $C\sqrt{(d_{1}+d_{2})KT\log(KT/\delta)}$ with probability at least $1-4\delta$. Here, $C>0$ is a universal constant and $T$ is assumed to be $\tilde{\Omega}(1)$.